OpenStudy (anonymous):

Reduce the following fraction: -32ABC/64AsqrdBCcubed a)-1/2ACsqrd b)-16AC/32ACsqrd c)1ACsqrd/2 d)-32/64ACsqrd

OpenStudy (campbell_st):

look at the problem in parts \[\frac{-32}{64} = \] \[\frac{A}{A^2} = \] \[\frac{B}{B} = \] \[\frac{C}{C^3}\] the solution is the product of the 4 answers above

OpenStudy (anonymous):

please help, im dumb lol

OpenStudy (campbell_st):

whats 32/64 =

OpenStudy (anonymous):

is tht the answer?

OpenStudy (campbell_st):

\[\frac{-32}{64} = \frac{-1}{2}\] \[\frac{A}{A^2} = \frac{1}{A}\] \[\frac{B}{B} = 1\] \[\frac{C}{C^3} = \frac{1}{C^2}\] so your answer is \[\frac{-1}{2} \times \frac{1}{A} \times \frac{1}{C^2} \times 1\] just write it correctly

OpenStudy (anonymous):

kay thanks